# A Statistical Machine Learning Approach for Adapting Reduced-Order Models using Projected Gaussian Process

### v1

@@ 2 Preliminaries on POD basis interpolation
This section provides preliminaries on the interpolation problem of POD basis modes. As discussed in Section 1.1, due to the computational cost associated with solving full-order governing equations, it is only possible to obtain the snapshot matrices for a relatively small number of parameters, $\bm{\Lambda}=\{\bm{\theta}_{1},\bm{\theta}_{2},\cdots,\bm{\theta}_{k}\}$, and obtain the optimal POD basis $\bm{\Phi}_{1},\bm{\Phi}_{2},\cdots,\bm{\Phi}_{k}$ for each parameter in $\bm{\Lambda}$. Hence, for any new parameter $\bm{\theta}^{*}\notin\bm{\Lambda}$, the interpolation of POD basis is needed to obtain $\bm{\Phi}^{*}$.
Grassmann Manifold. In honour of Hermann Grassmann, a Grassmann Manifold, $\mathcal{G}(r,n):=\{\mathbb{W}\subset\mathbb{R}^{n},\mathrm{dim}(\mathbb{W})=r\}$, is a differentiable manifold that collects all $r$-dimensional linear subspaces in $\mathbb{R}^{n}$[5]. Hence, the subspace $\mathbb{W}$ is a point in $\mathcal{G}(r,n)$ and the optimization problem in (2) essentially seeks the optimal point in $\mathcal{G}(r,n)$ that minimizes the distance $\sum_{i=1}^{n_{T}}||\bm{x}(t_{i};\bm{\theta})-\mathcal{W}(\bm{x}(t_{i};\bm{\theta}))||^{2}$.
Stiefel manifold and a submersion. A subspace in $\mathcal{G}(r,n)$ is spanned by an orthogonal basis $\bm{\Phi}\in\mathbb{R}^{n\times r}$ ($\bm{\Phi}^{T}\bm{\Phi}=\bm{I}_{r}$). A collection of $n\times r$ matrices with orthogonal columns forms a compact Stiefel manifold denoted by $\mathcal{ST}(r,n)$[12]. Note that, if $\bm{\Phi}$ spans a subspace $p\in\mathcal{G}(r,n)$, so does $\bm{\Phi}\bm{J}$ for any $r\times r$ matrix $\bm{J}$ such that $\bm{J}^{T}\bm{J}=\bm{I}_{p}$. Therefore, there exists is a fiber bundle
(3) $\pi:\bm{\Phi}\in\mathcal{ST}(r,n)\mapsto\pi(\bm{\Phi})=p\in\mathcal{G}(r,n)$
which indicates that any point $p\in\mathcal{G}(r,n)$ can be represented by a point of $\pi^{-1}(p)$.
@@
(4) $\left\langle v_{1},v_{2}\right\rangle:=\left\langle\bm{Z}_{1},\bm{Z}_{2}\right\rangle=\mathrm{trace}(\bm{Z}_{1}^{T}\bm{Z}_{2})$
where $\bm{Z}_{1}$ and $\bm{Z}_{2}$ are $n\times r$ matrices from the horizontal space of $\bm{\Phi}=\pi^{-1}(p)$ denoted by $\mathbb{H}_{\bm{\Phi}}:=\{\bm{Z}\in\mathbb{R}^{n\times r};\bm{Z}^{T}\bm{\Phi}=0\}$. In other words, given a base point $p$, the Grassmann Manifold inherits the Riemannian structure of the Stiefel manifold due to (3) and becomes a Riemannian manifold.
@@
(5) $\text{Exp}_{p}:v\in\mathcal{T}_{p}\mathcal{G}(r,n)\mapsto\pi(\bm{\Phi}\bm{V}_{\bm{Z}}\cos(\bm{\Sigma}_{\bm{Z}})+\bm{U}_{\bm{Z}}\sin(\bm{\Sigma}_{\bm{Z}}))\in\mathcal{G}(r,n)$
where $\bm{\Phi}=\pi^{-1}(p)$ and $\bm{Z}=\bm{U}_{\bm{Z}}\bm{\Sigma}_{\bm{Z}}\bm{V}_{\bm{Z}}$ is the SVD of $\bm{Z}$.
For any point $p\in\mathcal{G}(r,n)$ and a open set $U_{\bm{p}}:=\{p^{\prime}\in\mathcal{G}(r,n);\pi(\bm{\Phi})=p,\pi(\bm{\Phi}^{\prime})=p^{\prime},\pi(\bm{\Phi})^{T}\pi(\bm{\Phi}^{\prime})\text{ is invertible)}$), the Logarithm map is defined as
(6) $\text{Log}_{p}:p^{\prime}\in U_{\bm{p}}\mapsto\text{Log}_{p}(p^{\prime})\in\mathcal{T}_{p}\mathcal{G}(r,n).$
Existing POD interpolation method. Within $U_{\bm{p}}$, the Logarithm map $\text{Log}_{p}$ is a diffemorphism, and the one-one correspondence between $v\in\mathcal{T}_{p}\mathcal{G}$ and $p^{\prime}\in U_{\bm{p}}$ can be defined through the Exponential and Logarithm maps[12]. Hence, [2] described a well-known Lagrangian interpolation method for POD basis. In a nutshell, given the training data $\{\bm{\theta}_{i},\bm{\Phi}_{i}\}_{i=1}^{k}$, subspaces corresponding to $\bm{\theta}_{i}$ are mapped to vectors $v_{i}$ in the tangent space through the Logarithm map. Then, the interpolation of vectors can be done in the tangent space, and the interpolated vector $v^{*}$ is mapped back to the Grassmann Manifold through the Exponential map. Such a POD basis interpolation algorithm is provided in Algorithm 1.
Algorithm 1 POD basis interpolation on Grassmann Manifold
  Input: Integers $r$ and $n$ ($2r\leq n$), $\{\bm{\theta}_{i},\bm{\Phi}_{i}\}_{i=1}^{k}$, and a target parameter $\bm{\theta}^{*}$.
@@
    (7) $\bm{Z}_{i}=\bm{U}_{i}\text{arctan}(\bm{\Sigma}_{i})\bm{V}_{i}^{T}.$
@@
    (8) $\bm{Z}^{*}=\sum_{i=1}^{k}\prod_{i\neq j}\frac{\bm{\theta}^{*}-\bm{\theta}_{j}}{\bm{\theta}_{i}-\bm{\theta}_{j}}\bm{Z}_{i}=\bm{U}^{*}\bm{\Sigma}^{*}\bm{V}^{*}$
@@
    (9) $\bm{\Phi}^{*}=\bm{\Phi}_{0}\bm{V}^{*}\text{cos}\bm{\Sigma}^{*}+\bm{U}^{*}\text{sin}\bm{\Sigma}^{*}$
@@
  return  The interpolated POD basis $\bm{\Phi}^{*}$ for parameter $\bm{\theta}^{*}$.
3 Projected Gaussian Process for Adapting ROMs
In this section, we describe a supervised statistical learning method, based on the proposed pGP, for adapting ROMs against parameter changes.
3.1 Projection from $\mathbb{R}^{nr-r}$ to $\mathcal{G}(r,n)$
@@
From the preliminaries in Sec.2, there is an isomorphism between the horizontal space $\mathbb{H}_{\bm{\Phi}}$ and the tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$, i.e., $\pi_{\mathbb{H}}:\mathbb{H}_{\bm{\Phi}}\rightarrow\mathcal{T}_{p}\mathcal{G}(r,n)$. Hence, it is possible to find a unique $\bm{Z}\in\mathbb{H}_{\bm{\Phi}}$ such that $\pi_{\mathbb{H}}(\bm{Z})=v\in\mathcal{T}_{p}\mathcal{G}(r,n)$, and the matrix $\bm{Z}$ is called the horizontal lift of the vector $v\in\mathcal{T}_{p}\mathcal{G}(r,n)$. The existence of such an isomorphism may suggest that a matrix-variate GP can be employed to model the relationship between $\bm{\theta}$ and the matrix $\bm{Z}$. However, it is important to point out that this seemingly natural approach needs to be carefully carried out because the matrix $\bm{Z}$ should be defined on the horizontal space $\mathbb{H}_{\bm{\Phi}}$ given a base point $p$, rather than the Euclidean space $\mathbb{R}^{n\times r}$. Precisely speaking, if $\bm{Z}$ is modeled as a GP with respect to the parameter $\bm{\theta}$ on $\mathbb{R}^{n\times r}$, then, the point prediction, $\hat{\bm{Z}}^{*}$, is given by a linear combination of $\bm{Z}_{1},\bm{Z}_{2},\cdots,\bm{Z}_{k}$ based on the well-known results of GP regression[28]. Since $\bm{Z}_{1},\bm{Z}_{2},\cdots,\bm{Z}_{k}$ are in the horizontal space $\mathbb{H}_{\bm{\Phi}}$ , the point prediction $\hat{\bm{Z}}^{*}$ remains in the same space. However, the uncertainty quantification or predictive internal of $\hat{\bm{Z}}^{*}$ (or, the predicted $\hat{\bm{\Phi}}^{*}$) is problematic because a sample $\bm{Z}$ from a Gaussian distribution on $\mathbb{R}^{n\times r}$ does not necessarily belong to $\mathbb{H}_{\bm{\Phi}}$.
To remedy this issue, we first present the following results:
Proposition 3.1.
Given a basepoint $p\in\mathcal{G}(r,n)$ and $\bm{\Phi}=[\bm{\phi}_{1},\cdots,\bm{\phi}_{r}]$ such that $\pi(\bm{\Phi})=p$, let $\bm{y}$ be a $(nr-r)\times 1$ vector, $\tilde{\bm{\Phi}}=\mathrm{diag}\{\bm{\phi}_{1}^{T},\bm{\phi}_{2}^{T},\cdots,\bm{\phi}_{r}^{T}\}$ be a $r\times nr$ matrix with orthogonal rows, and $\bm{F}=[\bm{F}_{1}|\bm{F}_{2}|\cdots|\bm{F}_{r}]$ be a $(nr-r)\times nr$ matrix where each block $\bm{F}_{i}$, $i=1,2,\cdots,r$, is a $(nr-r)\times n$ matrix that satisfies the following conditions: (i) for any block $i$ ($i=1,2,\cdots,r$), only row $(i-1)\times(n-1)+1$ to row $i\times(n-1)$ contain non-zero entries; (ii) all non-zero row vectors are unit vectors; and (iii) all non-zero row vectors are orthogonal to each other (note that conditions (i)-(iii) imply $\tilde{\bm{\Phi}}\bm{F}^{T}=\bm{0}$). Then,
(10) $\bm{Z}=\mathrm{Mat}_{n,r}(\bm{F}^{T}\bm{y})$
is a $n\times r$ matrix that belongs to the horizontal space $\mathbb{H}_{\bm{\Phi}}$, i.e., $\bm{Z}^{T}\bm{\Phi}=\bm{0}$.
Proof. Note that $\bm{Z}$ belongs to the horizontal space of $\bm{\Phi}$, i.e.,
(11) $\tilde{\bm{\Phi}}\text{vec}(\bm{Z})=0$
@@
where $\text{vec}(\bm{Z})$ is the column-wise vectorization of $\bm{Z}$, and $\tilde{\bm{\Phi}}=\mathrm{diag}\{\bm{\phi}_{1}^{T},\bm{\phi}_{2}^{T},\cdots,\bm{\phi}_{r}^{T}\}$ is a $r\times nr$ block diagonal matrix with orthogonal rows. Then, by introducing a $(nr-r)\times nr$ matrix $\bm{F}$ with orthogonal rows such that $\tilde{\bm{\Phi}}\bm{F}^{T}=\bm{0}$, we have
(12) $\begin{pmatrix}\tilde{\bm{\Phi}}\\ \bm{F}\end{pmatrix}\text{vec}(\bm{Z})=\begin{pmatrix}\bm{0}\\ \bm{y}\end{pmatrix}$
where $\bm{y}$ is a $(nr-r)$-dimensional column vector.
Because $\bm{F}\text{vec}(\bm{Z})=\bm{y}$ is an underdetermined system, we may find $\text{vec}(\bm{Z})$ by $\bm{F}^{\dagger}\bm{y}$ where $\bm{F}^{\dagger}=\bm{F}^{T}(\bm{F}\bm{F}^{T})^{-1}$. From conditions (i)-(iii) above, $\bm{F}\bm{F}^{T}$ is an identity matrix, and $\bm{F}^{\dagger}\bm{F}^{T}$ as shown in Proposition 3.1.
Here, the choice of $\bm{F}$ depends on $\bm{\Phi}$ (and the basepoint $\bm{p}$ in Proposition 3.1) as it must satisfy conditions (i)-(iii). Let $\tilde{\bm{\Phi}}=[\tilde{\bm{\Phi}}_{1},\tilde{\bm{\Phi}}_{2},\cdots,\tilde{\bm{\Phi}}_{r}]$, and let $\tilde{\bm{F}}=[\tilde{\bm{F}}_{1},\tilde{\bm{F}}_{2},\cdots,\tilde{\bm{F}}_{r}]$ be a $(nr-r)\times nr$ proposal matrix where each block $\tilde{\bm{F}}_{i}$ has the structure as specified by condition (i) above. Then, applying the Gram-Schmidt (G-S) process can be applied to the row vectors of each block $\bigl{(}\begin{smallmatrix}\tilde{\bm{\Phi}}_{i}\\ \tilde{\bm{F}}_{i}\end{smallmatrix}\bigr{)}$ (for $i=1,2,\cdots,r$) and normalizing each non-zero rows after the G-S process, we obtain $\bigl{(}\begin{smallmatrix}\tilde{\bm{\Phi}}_{i}\\ \bm{F}_{i}\end{smallmatrix}\bigr{)}$ where $\bm{F}_{i}$ satisfies conditions (i)-(iii). Note that, the G-S process starts from the $i$-th row because the leading $(i-1)$ rows of $\tilde{\bm{\Phi}}_{i}$ are zeros.
@@
Proposition 3.1 implies that, from a vector $\bm{y}\in\mathbb{R}^{nr-r}$, it is possible to compute a matrix $\bm{Z}=\mathrm{Mat}_{n,r}(\bm{F}^{\dagger}\bm{y})$ that comes from the horizontal space $\mathbb{H}_{\bm{\Phi}}$. In fact, it is not hard to understand why the dimension of $\bm{y}$ is $nr-r$. Although there are $n\times r$ elements in the matrix $\bm{Z}$, the constraint $\bm{Z}^{T}\bm{\Phi}=\bm{0}$ (i.e., $\bm{Z}$ must come from the horizontal space of $\bm{\Phi}$) takes $r$ degree of freedom away, leaving behind $nr-r$ degrees of freedom. Hence, for any basepoint $p\in\mathcal{G}(r,n)$ and the tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$, we can define the following mapping from $\mathbb{R}^{nr-r}$ to $\mathcal{G}(r,n)$,
(13) $\mathcal{P}_{p}:\bm{y}\in\mathbb{R}^{nr-r}\mapsto\pi(\bm{\Phi}\bm{V}_{\bm{Z}}\cos(\bm{\Sigma}_{\bm{Z}})+\bm{U}_{\bm{Z}}\sin(\bm{\Sigma}_{\bm{Z}}))\in\mathcal{G}(r,n).$
@@
Here, $\mathcal{P}_{p}:=\tilde{\mathcal{P}}_{p}\circ\text{Exp}_{p}$ where $\tilde{\mathcal{P}}_{p}$ maps an $(nr-r)$-dimensional vector to the tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$,
(14) $\tilde{\mathcal{P}}_{p}:\bm{y}\in\mathbb{R}^{nr-r}\mapsto v\in\mathcal{T}_{p}\mathcal{G}(r,n)$

### v2

@@ 2 Preliminaries on POD basis interpolation
This section provides preliminaries on the interpolation of POD bases. Suppose that the snapshot data matrices are generated at a set of parameter settings, $\bm{\Lambda}=\{\bm{\theta}_{1},\bm{\theta}_{2},\cdots,\bm{\theta}_{k}\}$, and the optimal POD bases $\bm{\Phi}_{1},\bm{\Phi}_{2},\cdots,\bm{\Phi}_{k}$ are obtained for each parameter in $\bm{\Lambda}$. Then, the POD basis interpolation problem involves finding the POD basis $\bm{\Phi}^{*}$ at new parameters $\bm{\theta}^{*}\notin\bm{\Lambda}$. Some important concepts are firstly introduced.
[…]
Stiefel manifold and a submersion. A subspace in $\mathcal{G}(r,n)$ is spanned by an orthogonal basis $\bm{\Phi}\in\mathbb{R}^{n\times r}$ ($\bm{\Phi}^{T}\bm{\Phi}=\bm{I}_{r}$). A collection of $n\times r$ matrices with orthogonal columns forms a compact Stiefel manifold denoted by $\mathcal{ST}(r,n)$. Note that, if $\bm{\Phi}$ spans a subspace $p\in\mathcal{G}(r,n)$, so does $\bm{\Phi}\bm{J}$ for any $r\times r$ matrix $\bm{J}$ such that $\bm{J}^{T}\bm{J}=\bm{I}_{p}$. Therefore, there exists a fiber bundle
$\pi:\bm{\Phi}\in\mathcal{ST}(r,n)\mapsto\pi(\bm{\Phi})=p\in\mathcal{G}(r,n)$ (3)
which indicates that any point $p\in\mathcal{G}(r,n)$ can be represented by a point of the fiber $\pi^{-1}(p)$ (Friderikos et al.,, 2020).
@@
$\left\langle v_{1},v_{2}\right\rangle:=\left\langle\bm{Z}_{1},\bm{Z}_{2}\right\rangle=\mathrm{trace}(\bm{Z}_{1}^{T}\bm{Z}_{2})$ (4)
where $\bm{Z}_{1}$ and $\bm{Z}_{2}$ are $n\times r$ matrices from the horizontal space of $\bm{\Phi}\in\pi^{-1}(p)$ denoted by $\mathbb{H}_{\bm{\Phi}}:=\{\bm{Z}\in\mathbb{R}^{n\times r};\bm{Z}^{T}\bm{\Phi}=0\}$. In other words, given a base point $p$, the Grassmann manifold inherits the Riemannian structure of the Stiefel manifold due to (3) and is a Riemannian manifold.
@@
$\text{Exp}_{p}:v\in\mathcal{T}_{p}\mathcal{G}(r,n)\mapsto\pi(\bm{\Phi}\bm{V}_{\bm{Z}}\cos(\bm{\Sigma}_{\bm{Z}})+\bm{U}_{\bm{Z}}\sin(\bm{\Sigma}_{\bm{Z}}))\in\mathcal{G}(r,n)$ (5)
where $\bm{\Phi}\in\pi^{-1}(p)$ and $\bm{Z}=\bm{U}_{\bm{Z}}\bm{\Sigma}_{\bm{Z}}\bm{V}_{\bm{Z}}$ is the SVD of $\bm{Z}\in\mathbb{H}_{\bm{\Phi}}$. Note that, for a given $\Phi$ such that $\pi(\Phi)=p$, the tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$ is isomorphic to the horizontal space $\mathbb{H}_{\bm{\Phi}}$. For any $v\in\mathcal{T}_{p}\mathcal{G}(r,n)$, there is a unique $\bm{Z}\in\mathbb{H}_{\bm{\Phi}}$ known as the horizontal lift of $v$ (Friderikos et al.,, 2020).
For any point $p\in\mathcal{G}(r,n)$ and a open set $U_{\bm{p}}:=\{p^{\prime}\in\mathcal{G}(r,n)\}$, the Logarithm map is defined as
$\text{Log}_{p}:p^{\prime}\in U_{\bm{p}}\mapsto\text{Log}_{p}(p^{\prime})\in\mathcal{T}_{p}\mathcal{G}(r,n).$ (6)
Existing POD interpolation method. Within $U_{\bm{p}}$, the Logarithm map $\text{Log}_{p}$ is a diffemorphism, and the one-to-one correspondence between $v\in\mathcal{T}_{p}\mathcal{G}$ and $p^{\prime}\in U_{\bm{p}}$ can be defined through the Exponential and Logarithm maps (Friderikos et al.,, 2020). Hence, Amsallem and Farhat, (2008) described a popular Lagrangian interpolation method for POD bases. In a nutshell, given the training data $\{\bm{\theta}_{i},\bm{\Phi}_{i}\}_{i=1}^{k}$, subspaces corresponding to $\bm{\theta}_{i}$ are mapped to vectors $v_{i}$ in the tangent space through the Logarithm map. Then, the interpolation of vectors can be done in the tangent space, and the interpolated vector $v^{*}$ is mapped back to the Grassmann manifold through the Exponential map. Algorithm 2.1 provides an example of such an approach for the one-dimensional case, i.e., $\theta\in\mathbb{R}$.
 Input: Integers $r$ and $n$ ($2r\leq n$), $\{\theta_{i},\bm{\Phi}_{i}\}_{i=1}^{k}$, and a target parameter $\theta^{*}$.
@@
    $\bm{Z}_{i}=\bm{U}_{i}\text{arctan}(\bm{\Sigma}_{i})\bm{V}_{i}^{T}.$ (7)
@@
    $\bm{Z}^{*}=\sum_{i=1}^{k}\prod_{i\neq j}\frac{\theta^{*}-\theta_{j}}{\theta_{i}-\theta_{j}}\bm{Z}_{i}=\bm{U}^{*}\bm{\Sigma}^{*}\bm{V}^{*}.$ (8)
    Note that, when the dimension of the parameter space $d$ is larger than one, the Lagrangian interpolation in (8) can be replaced by other multidimensional interpolation methods.
@@
    $\bm{\Phi}^{*}=\bm{\Phi}_{0}\bm{V}^{*}\text{cos}\bm{\Sigma}^{*}+\bm{U}^{*}\text{sin}\bm{\Sigma}^{*}$ (9)
@@
return  The interpolated POD basis $\bm{\Phi}^{*}$ for parameter $\theta^{*}$.
Algorithm 1 POD basis interpolation on Grassmann manifold
3 Projected Gaussian Process for Adapting ROM
In this section, we present a supervised statistical learning method, based on the proposed pGP, for adapting POD bases against parameters change.
3.1 A projection from $\mathbb{R}^{nr-r}$ to $\mathcal{G}(r,n)$
@@
From the preliminaries in Sec.2, there is an isomorphism between the horizontal space $\mathbb{H}_{\bm{\Phi}}$ and the tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$, i.e., $\pi_{\mathbb{H}}:\mathbb{H}_{\bm{\Phi}}\rightarrow\mathcal{T}_{p}\mathcal{G}(r,n)$. Hence, it is possible to find a unique $\bm{Z}\in\mathbb{H}_{\bm{\Phi}}$ such that $\pi_{\mathbb{H}}(\bm{Z})=v\in\mathcal{T}_{p}\mathcal{G}(r,n)$, and the matrix $\bm{Z}$ is called the horizontal lift of the vector $v\in\mathcal{T}_{p}\mathcal{G}(r,n)$. The existence of such an isomorphism may suggest that a matrix-variate GP can be employed to model the relationship between $\bm{\theta}$ and the matrix $\bm{Z}$. However, it is important to point out that this seemingly natural approach needs to be carefully carried out because the matrix $\bm{Z}$ should be defined on the horizontal space $\mathbb{H}_{\bm{\Phi}}$ given a base point $p$, rather than the Euclidean space $\mathbb{R}^{n\times r}$. Precisely speaking, if $\bm{Z}$ is modeled as a GP with respect to the parameter $\bm{\theta}$ on $\mathbb{R}^{n\times r}$, then, the point prediction, $\hat{\bm{Z}}^{*}$, is given by a linear combination of $\bm{Z}_{1},\bm{Z}_{2},\cdots,\bm{Z}_{k}$ based on the well-known results of GP regression (Rasmussen and Williams,, 2005). Since $\bm{Z}_{1},\bm{Z}_{2},\cdots,\bm{Z}_{k}$ are in the horizontal space $\mathbb{H}_{\bm{\Phi}}$ , the point prediction $\hat{\bm{Z}}^{*}$ remains in the same space. However, the uncertainty quantification or predictive internal of $\hat{\bm{Z}}^{*}$ (or, the predicted $\hat{\bm{\Phi}}^{*}$) is problematic because a sample $\bm{Z}$ from a Gaussian distribution on $\mathbb{R}^{n\times r}$ does not necessarily belong to $\mathbb{H}_{\bm{\Phi}}$.
To remedy this issue, we leverage the following result:
Proposition 3.1. For a given basepoint $p\in\mathcal{G}(r,n)$ and its corresponding basis matrix $\bm{\Phi}=[\bm{\phi}_{1},\cdots,\bm{\phi}_{r}]$, i.e, $\pi(\bm{\Phi})=p$, it is possible to obtain a $n\times r$ matrix $\bm{Z}$ from a $(nr-r)\times nr$ matrix $\bm{F}$ and a $(nr-r)\times 1$ vector $\bm{y}$ as follows
$\bm{Z}=\mathrm{Mat}_{n,r}(\bm{F}^{T}\bm{y})$ (10)
@@
such that $\bm{Z}$ belongs to the horizontal space $\mathbb{H}_{\bm{\Phi}}$, i.e., $\bm{Z}^{T}\bm{\Phi}=\bm{0}$. Here, the matrix $\bm{F}$ has orthogonal rows and satisfies the conditions $\tilde{\bm{\Phi}}\bm{F}^{T}=\bm{0}$ and $\bm{F}\bm{F}^{T}=\bm{I}$, where $\tilde{\bm{\Phi}}=\mathrm{diag}\{\bm{\phi}_{1}^{T},\bm{\phi}_{2}^{T},\cdots,\bm{\phi}_{r}^{T}\}$ is a $r\times nr$ matrix with orthogonal rows.
To prove Proposition 3.1, note that,
$\tilde{\bm{\Phi}}\text{vec}(\bm{Z})=0$ (11)
when $\bm{Z}$ belongs to the horizontal space of $\bm{\Phi}$, i.e., $\bm{Z}^{T}\bm{\Phi}=\bm{0}$. Then, if we can choose a $(nr-r)\times nr$ matrix $\bm{F}$ which has orthogonal rows and satisfies $\tilde{\bm{\Phi}}\bm{F}^{T}=\bm{0}$, we have
$\begin{pmatrix}\tilde{\bm{\Phi}}\\ \bm{F}\end{pmatrix}\text{vec}(\bm{Z})=\begin{pmatrix}\bm{0}\\ \bm{y}\end{pmatrix}$ (12)
where $\bm{y}$ is a (nrr)(nr-r)-dimensional column vector. Because $\bm{F}\text{vec}(\bm{Z})=\bm{y}$ is an underdetermined system, one can find $\text{vec}(\bm{Z})$ by $\bm{F}^{\dagger}\bm{y}$ where $\bm{F}^{\dagger}=\bm{F}^{T}(\bm{F}\bm{F}^{T})^{-1}$. Next, if we further impose the condition $\bm{F}\bm{F}^{T}=\bm{I}$, then, $\bm{F}^{\dagger}=\bm{F}^{T}$ and $\text{vec}(\bm{Z})=\bm{F}^{T}\bm{y}$ as shown in Proposition 3.1. It is straightforward to verify that, when $\text{vec}(\bm{Z})=\bm{F}^{T}\bm{y}$, we have $\tilde{\bm{\Phi}}\text{vec}(\bm{Z})=\tilde{\bm{\Phi}}\bm{F}^{T}\bm{y}=\bm{0}$, which implies that $\bm{Z}$ belongs to the horizontal space of $\bm{\Phi}$ according to (11).
@@
Proposition 3.1 implies that, from a vector $\bm{y}\in\mathbb{R}^{nr-r}$, it is possible to compute a matrix $\bm{Z}=\mathrm{Mat}_{n,r}(\bm{F}^{\dagger}\bm{y})$ that comes from the horizontal space $\mathbb{H}_{\bm{\Phi}}$. In fact, it is not hard to understand why the dimension of $\bm{y}$ is nrrnr-r. Although there are $n\times r$ elements in the matrix $\bm{Z}$, the constraint $\bm{Z}^{T}\bm{\Phi}=\bm{0}$ consists of a number of $r$ linear equations (i.e., $\bm{Z}$ must come from the horizontal space of $\bm{\Phi}$) takes away $r$ degree of freedom, leaving behind nrrnr-r degrees of freedom.
It is also noted that, the choice of $\bm{F}$ is obviously not unique as long as the matrix has orthogonal columns and satisfies $\tilde{\bm{\Phi}}\bm{F}^{T}=\bm{0}$ and $\bm{F}\bm{F}^{T}=\bm{I}$. In our implementation, a computationally efficient procedure is used to construct the matrix $\bm{F}$. We let $\tilde{\bm{\Phi}}=[\tilde{\bm{\Phi}}_{1},\tilde{\bm{\Phi}}_{2},\cdots,\tilde{\bm{\Phi}}_{r}]$, and let $\tilde{\bm{F}}=[\tilde{\bm{F}}_{1},\tilde{\bm{F}}_{2},\cdots,\tilde{\bm{F}}_{r}]$ be a $(nr-r)\times nr$ proposal matrix where each block $\tilde{\bm{F}}_{i}$ contains non-zero entries only from row $(i-1)\times(n-1)+1$ to row $i\times(n-1)$. Then, the Gram-Schmidt (G-S) process can be applied, in parallel, to the row vectors of each block $\bigl{(}\begin{smallmatrix}\tilde{\bm{\Phi}}_{i}\\ \tilde{\bm{F}}_{i}\end{smallmatrix}\bigr{)}$ for $i=1,2,\cdots,r$. After the G-S process, each non-zero row is normalized, yielding $\bigl{(}\begin{smallmatrix}\tilde{\bm{\Phi}}_{i}\\ \bm{F}_{i}\end{smallmatrix}\bigr{)}$ and $\bm{F}=[\bm{F}_{1},\bm{F}_{2},\cdots,\bm{F}_{r}]$ that satisfies the conditions needed, i.e., orthogonal rows, $\tilde{\bm{\Phi}}\bm{F}^{T}=\bm{0}$ and $\bm{F}\bm{F}^{T}=\bm{I}$.
Following Proposition 3.1, for any basepoint $p\in\mathcal{G}(r,n)$ and the tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$, we can define the following mapping from $\mathbb{R}^{nr-r}$ to $\mathcal{G}(r,n)$,
$\mathcal{P}_{p}:\bm{y}\in\mathbb{R}^{nr-r}\mapsto\pi(\bm{\Phi}\bm{V}_{\bm{Z}}\cos(\bm{\Sigma}_{\bm{Z}})+\bm{U}_{\bm{Z}}\sin(\bm{\Sigma}_{\bm{Z}}))\in\mathcal{G}(r,n).$ (13)
@@
Here, $\mathcal{P}_{p}:=\tilde{\mathcal{P}}_{p}\circ\text{Exp}_{p}$ where $\tilde{\mathcal{P}}_{p}$ maps an (nrr)(nr-r)-dimensional vector to the tangent space $\mathcal{T}_{p}\mathcal{G}(r,n)$,
$\tilde{\mathcal{P}}_{p}:\bm{y}\in\mathbb{R}^{nr-r}\mapsto v\in\mathcal{T}_{p}\mathcal{G}(r,n)$ (14)